# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502 Japan

### v2

@@ 6 Correlation functions and transverse suppression
Let us compare this result with the field theory shown in Section 2. As explained, the two-point function (7) is proportional to $\quantity(\delta_{ij}-\alpha\frac{x_{i}x_{j}}{x^{2}})$. If we take $\vec{x}=(x,0,0)$, then the second term has a spin-component dependence, resulting in the total ratio of $(1-\alpha)$ for $i=j=x$ and $1$ for $i=j=y$ ( or $z$). For small $\eta$, $\alpha$ is negative, so the field theory predicts the suppression of the transverse direction (i.e. $yy$ and $zz$) compared with the longitudinal one (i.e. $xx$). This is indeed seen in our numerical plots.
@@
Next, we estimate the critical exponent related to the anomalous dimension from the correlation functions. We fit the diagonal elements of $C_{ij}(\vec{x})$, assuming a power function $c/|\vec{x}|^{p}$ in long $|\vec{x}|$ regimes. Since the diagonal elements are symmetric under $\vec{x}\to-\vec{x}$, we take the forward-backward average, $[C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$, before fitting. The fitting range is chosen as a region where the power-law behavior is observed. For $\lambda=0$, we set the range to $|\vec{x}|\in[6,16]$ for all the diagonal elements. As for $\lambda=4$ and $8$, the range is $|\vec{x}|\in[7,16]$ for the $xx$ element and $|\vec{x}|\in[4,18]$ for the others. The fitting results are summarized in Table 3 and plotted in Figure 14 with the data of the forward-backward average.
@@
$xx$ 1.02(4) 0.0910(67) 0.31 $[6,16]$
0 $yy$ 0.93(4) 0.0741(54) 0.28 $[6,16]$
$zz$ 1.01(4) 0.0872(71) 0.94 $[6,16]$
$xx$ 0.60(5) 0.0223(22) 1.61 $[7,16]$
4 $yy$ 0.35(3) 0.0114(6) 1.01 $[4,18]$
$zz$ 0.31(3) 0.0103(6) 0.95 $[4,18]$
$xx$ 0.60(5) 0.0208(23) 0.98 $[7,16]$
8 $yy$ 0.30(2) 0.0097(5) 0.87 $[4,18]$
$zz$ 0.31(3) 0.0096(5) 0.51 $[4,18]$
@@
Figure 14: The forward-backward averages $[C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$ of the diagonal elements of the correlation function are plotted against $|\vec{x}|$ in log-log scale. The left, center, and right panels show the $xx$, $yy$, and $zz$ elements, where the circle-blue, square-green, and diamond-red symbols correspond to $\lambda=0$, $4$, and $8$, respectively. The solid lines depict the fitting results by $c/|\vec{x}|^{p}$.
@@
The power $p=2\Delta_{\phi}$ is related to the critical exponent $\eta$ via Eq. (11) as $\Delta_{\phi}=(1+\eta)/2$ with $d=3$. Indeed, we obtain $\Delta_{\phi}\approx 1/2$ from the correlation function for $\lambda=0$, which is consistent with $\eta=0.0318(27)(^{+10}_{-7})$ in Eq. (38) from the magnetic susceptibility as it is in the same order as the fitting error of $p$. As for $\lambda=8$, we have $\Delta_{\phi}\approx 0.3$ from the $xx$ component, [^5] We could have read $\Delta_{\phi}$ from $yy$ and $zz$ component, which gives different values here. At two-loop, the anomalous dimensions of $xx$ and $yy$ (or $zz$) components are different for finite $\lambda$ (the difference becomes zero in the $\lambda\to\infty$ limit). This may be a potential cause of the difference observed here. which implies a negative value of $\eta$. However, we obtained the positive value $\eta=0.132(8)(^{+0}_{-4})$ in Eq. (39) from the magnetic susceptibility. Thus, the two results obtained from the different observables appear inconsistent in the case of the local Heisenberg-dipolar model. Note that, in reflection positive theory such as the Heisenberg model, it is strictly proved that $\eta$ must be positive. On the other hand, this restriction does not apply to the local Heisenberg-dipolar model since it violates the reflection positivity. In addition, even in the continuum theory, it is known that the derivative interaction can cause a negative anomalous dimension for the scalar field theory Higashijima:2003et . We cannot exclude the possibility of the negative $\eta$ at the dipolar fixed point.
@@
Let us, however, comment on subtleties in the analysis of the correlation function we have observed. The power-law scaling region for $\lambda=8$ is relatively unclear than the case of $\lambda=0$. It is possible that the finite-size effect on the correlation function becomes more significant for $\lambda=8$ and affects the power-law behavior. Indeed, we did see some hints of this behavior by comparing the results for $L=32$ and $40$ in Appendix E. There, we see that the power $p$ read from the correlation functions of different lattice sizes differ: the larger the lattice size $L$, we had the larger $p$. It is possible that larger $L$ gives a consistent estimate of $\eta$ with the other method. We also observed that the correlation function is sensitive to the temperature, and thus it could be the source of ambiguity as well. Here we choose the peak temperature $\beta_{\mathrm{peak}}$ of the magnetic susceptibility, but the behavior of the correlation function changes by even slightly shifting the temperature. We examine the temperature dependence in Appendix F.
@@
All in all, we find reading critical exponents from the correlation function is less accurate than reading them from the finite size corrections of Binder parameters and magnetic susceptibility, both in the Heisenberg model and the local Heisenberg-dipolar model. This is in accord with what has been observed in the literature. We therefore report the values in Section 5 as our main results.
7 Summary and discussion
@@
Our obtained value of the critical exponent $\gamma$ turned out to be smaller than the estimation of the other method. There might exist a theoretical, as well as experimental, interpretation of this apparent smallness (if we assumed that it is indeed different from the dipolar fixed point value $\gamma_{\mathrm{dipolar}}$). [^6] The following interpretation was communicated to us by Slava Rychkov. We would like to thank him for his insight. Recall that our simulation at finite $\lambda$ does not correspond to a true RG fixed point but rather to a crossover point. In such cases, it is argued that the effective susceptibility exponent $\gamma_{\mathrm{eff}}=\frac{d\log\chi}{d\log\beta}$ may show a significant dip away from the fixed point Gimenez-Grau:2023lpz . In our case, the deviation from the RG fixed point is mainly caused by finite $\lambda$, and our definition of $\gamma$ is different from $\gamma_{\mathrm{eff}}$, but it is plausible that the apparent smallness of our obtained $\gamma$ may be attributed to the similar effects. We also note that there is experimental evidence for such an effect in dipolar ferromagnets (see e.g. inbook and reference therein).
A future direction is the improvement of simulation algorithms. In this paper, we used the simplest Metropolis algorithm and encountered severe autocorrelation around the critical temperature, namely the critical slowing down, which prevented us from taking a larger lattice size than $L=40$, in particular with larger $\lambda$. For more precise determination of the critical exponents, the cluster algorithm PhysRevLett.58.86 ; PhysRevLett.62.361 can deal with this problem since it efficiently reduces the autocorrelation by a non-local update of the spin configuration. However, it is not straightforward to apply the algorithm to the local Heisenberg-dipolar model due to the next-nearest-neighbor interaction, the first term in the second line of Eq. (16). In this case, the generalized versions of the cluster algorithm will be helpful PhysRevLett.65.941 ; PhysRevB.43.8539 . Alternatively, it is also interesting to simulate the long-range dipolar interaction directly without using the local description. Indeed, the efficient algorithms for spin models with non-local interaction are proposed in Refs. doi:10.1142/S0129183195000265 ; Fukui_2009
@@
Finally, we wish there would be a continuous effort to experimentally verify the dipolar effects in real ferromagnets. [^7] In the Chinese character, the root of magnetism is associated with a stone with benevolence. Why? In ancient Chinese stories, it is said that the way the ferromagnet attracts iron rocks makes it look as if an infant is naturally attracted to his or her mother. Probably the dipolar magnet will embody more benevolence because the magnetism from it is divergenceless, so it appears to be more gentle and embracing. So far in human history, the distinct features of dipolar magnetism have been observed in esoteric materials such as Europium compounds, but the universal nature of the benevolence should also be seen in more conventional materials such as iron or nickel in future experiments, just with more care.
@@ Appendix B Hubbard-Stratonovich transformation
Mapping of the parameters in the continuum field theory of Eq. (9) and a lattice model in Section 3 is non-trivial. Specifically, in Eq. (9), the term that imposes an important constraint in the local Heisenberg-dipolar model may acquire a multiplicative factor when mapped to the lattice model at finite $\lambda$. We derive the factor from the Hubbard-Stratonovich transformation within the mean field approximation.
@@
$1=\quantity(\frac{\beta}{2\pi})^{|\Lambda|/2}\frac{1}{\sqrt{\det K}}\int\quantity(\prod_{l\in\Lambda}d\phi_{l})\exp\quantity(-\frac{\beta}{2}\sum_{n,m\in\Lambda}\vec{\phi}_{n}\cdot K^{-1}_{nm}\vec{\phi}_{m}),$ (46)
@@
[…]
@@
$Z\sim\int\mathcal{D}\vec{\phi}\exp\quantity(-\frac{\beta}{2}\int_{x,y}\vec{\phi}(x)\cdot K^{-1}(x-y)\vec{\phi}(y)+\frac{\pi\beta^{2}\mu^{2}}{3}\int_{x}(\vec{\phi}(x)+\vec{h})^{2}),$ (49)
@@
where $\hat{i}\in\{\hat{1},\hat{2},\hat{3}\}$ denotes the lattice unit vector in the $i$-th direction. We have added the last term (with $\theta>6$) as a trick because we want to make the matrix $K$ positive.
@@
$\int d^{3}xd^{3}y\ \vec{\phi}(x)\cdot\tilde{K}(x-y)\vec{\phi}(y)=\int d^{3}x\quantity(-J\theta\phi_{i}\phi_{i}+Ja^{2}(\partial_{i}\phi_{j})^{2}-2a^{2}\lambda(\partial_{i}\phi_{i})^{2}),$ (51)
@@
$\int d^{3}xd^{3}y\ \vec{\phi}(x)K^{-1}(x-y)\vec{\phi}(y)\simeq\int d^{3}x\quantity(-\frac{1}{J\theta}\phi_{i}\phi_{i}-\frac{a^{2}}{J\theta^{2}}(\partial_{i}\phi_{j})^{2}+\frac{2a^{2}\lambda}{J^{2}\theta^{2}}(\partial_{i}\phi_{i})^{2}).$ (52)
@@
Consequently, we show that the normalization of $\lambda$ here is the one introduced in section 2 when we take $J=-1$ in the lattice Hamiltonian (14) as a corresponding model.
@@ Appendix C Determination of the peak position of $\chi_{m}$
$\beta_{\mathrm{peak}}\,(L=32)$ 0.689135(17) 0.353046(8) 0.320598(12)
$\beta_{\mathrm{peak}}\,(L=40)$ 0.690190(12) 0.352883(11) 0.320374(6)
$\beta_{c}$ 0.693035(11) -- 0.319844(18)
@@ Appendix D Fitting results of the interpolating functions
Here, we summarize the best-fit values of the parameters $\{c_{i}(L)\}$ and $\{d_{i}(L)\}$ of the interpolating functions (30) and (36) which are obtained in Sections 5.2 and 5.5 in Table 5. The fitting errors of these values are large in some cases because the $\chi^{2}$-fit suffers from a relatively flat direction in the parameter spaces of $\{c_{i}(L)\}$ and $\{d_{i}(L)\}$. Even though the fitting parameters have large errors, the interpolated values have smaller errors, used to obtain the intersection and gradient, etc. Thus, it does not cause a problem in our analyses.
@@
Table 5: The best-fit values of the parameters $\{c_{i}(L)\}$ and $\{d_{i}(L)\}$ in the interpolating function (30) and (36) for $\lambda=0$ (top) and $8$ (bottom). These values are normalized by $L^{2}$ or $L^{4}{10}^{3}$ to be almost the same order in these tables.
Appendix E Finite size effect on the correlation function
Here, we discuss the system-size dependence of the correlation function. In Figure 16, the correlation functions at the different system sizes $L=32$ and $40$ are compared, where the forward-backward average $[C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$ is taken. The square-red data denote the result at the peak temperatures $\beta_{\mathrm{peak}}$ of the magnetic susceptibility on $L=40$ lattice, previously shown in Section 6, The circle-blue data are on $L=32$ lattice, where we tune the value of $\beta$ as $\beta_{\mathrm{peak}}=0.689135$ and $0.320598$ for $\lambda=0$ and $8$, respectively. We also depict the fitting lines of $c/|\vec{x}|^{p}$ in Figure 16, where the best-fit values are shown in Table 6 and 3 for $L=32$ and $40$, respectively.
Figure 16: The forward-backward averages $[C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$ of the diagonal elements of the correlation function are plotted against $|\vec{x}|$ in log-log scale for $\lambda=0$ (top) and $8$ (bottom). The left, center, and right columns show the $xx$, $yy$, and $zz$ elements, where the circle-blue and diamond-red symbols correspond to $L=32$ and $40$, respectively. The solid lines depict the fitting results by $c/|\vec{x}|^{p}$.

### v3

@@ VI Correlation functions and transverse suppression
Let us compare this result with the field theory shown in Section II. As explained, the two-point function (7) is proportional to $\quantity(\delta_{ij}-\alpha\frac{x_{i}x_{j}}{x^{2}})$. If we take $\vec{x}=(x,0,0)$, then the second term has a spin-component dependence, resulting in the total ratio of $(1-\alpha)$ for $i=j=x$ and $1$ for $i=j=y$ ( or $z$). For small $\eta$, $\alpha$ is negative, so the field theory predicts the suppression of the transverse direction (i.e. $yy$ and $zz$) compared with the longitudinal one (i.e. $xx$). This is indeed seen in our numerical plots.
@@
Next, we estimate the critical exponent related to the anomalous dimension from the correlation functions. We fit the diagonal elements of $C_{ij}(\vec{x})$, assuming a power function $c/|\vec{x}|^{p}$ in long $|\vec{x}|$ regimes. Since the diagonal elements are symmetric under $\vec{x}\to-\vec{x}$, we take the forward-backward average, $C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$, before fitting. The fitting range is chosen as a region where the power-law behavior is observed. For $\lambda=0$, we set the range to $|\vec{x}|\in[6,16]$ for all the diagonal elements. As for $\lambda=4$ and $8$, the range is $|\vec{x}|\in[7,16]$ for the $xx$ element and $|\vec{x}|\in[4,18]$ for the others. The fitting results are summarized in Table 3 and plotted in Figure 14 with the data of the forward-backward average.
@@
xxxx 1.02(4) 0.0910(67) 0.31 $[6,16]$
0 yyyy 0.93(4) 0.0741(54) 0.28 $[6,16]$
zzzz 1.01(4) 0.0872(71) 0.94 $[6,16]$
xxxx 0.60(5) 0.0223(22) 1.61 $[7,16]$
4 yyyy 0.35(3) 0.0114(6) 1.01 $[4,18]$
zzzz 0.31(3) 0.0103(6) 0.95 $[4,18]$
xxxx 0.60(5) 0.0208(23) 0.98 $[7,16]$
8 yyyy 0.30(2) 0.0097(5) 0.87 $[4,18]$
zzzz 0.31(3) 0.0096(5) 0.51 $[4,18]$
@@
Figure 14: The forward-backward averages $C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$ of the diagonal elements of the correlation function are plotted against $|\vec{x}|$ in log-log scale. The left, center, and right panels show the $xx$, yyyy, and $zz$ elements, where the circle-blue, square-green, and diamond-red symbols correspond to $\lambda=0$, $4$, and $8$, respectively. The solid lines depict the fitting results by $c/|\vec{x}|^{p}$.
@@
The power $p=2\Delta_{\phi}$ is related to the critical exponent $\eta$ via Eq. (11) as $\Delta_{\phi}=(1+\eta)/2$ with $d=3$. Indeed, we obtain $\Delta_{\phi}\approx 1/2$ from the correlation function for $\lambda=0$, which is consistent with $\eta=0.0318(27)(^{+10}_{-7})(^{+27}_{-28})_{\beta_{c}}$ in Eq. (38) from the magnetic susceptibility as it is in the same order as the fitting error of $p$. As for $\lambda=8$, we have $\Delta_{\phi}\approx 0.3$ from the $xx$ component,555 We could have read $\Delta_{\phi}$ from $yy$ and $zz$ component, which gives different values here. At two-loop, the anomalous dimensions of $xx$ and $yy$ (or $zz$) components are different for finite $\lambda$ (the difference becomes zero in the $\lambda\to\infty$ limit). This may be a potential cause of the difference observed here. which implies a negative value of $\eta$. However, we obtained the positive value $\eta=0.132(8)(^{+0}_{-4})(^{+12}_{-11})_{\beta_{c}}$ in Eq. (39) from the magnetic susceptibility. Thus, the two results obtained from the different observables appear inconsistent in the case of the local Heisenberg-dipolar model. Note that, in reflection positive theory such as the Heisenberg model, it is strictly proved that $\eta$ must be positive. On the other hand, this restriction does not apply to the local Heisenberg-dipolar model since it violates the reflection positivity. In addition, even in the continuum theory, it is known that the derivative interaction can cause a negative anomalous dimension for the scalar field theory [33]. We cannot exclude the possibility of the negative $\eta$ at the dipolar fixed point.
@@
Let us make some remarks on this analysis. First, to see the dipolar fixed point we have to consider the extrapolated value of $\eta$ in the $\lambda\to\infty$ limit. See Appendix  E for a detailed analysis. The next comments are subtleties in the analysis of the correlation function we have observed. The power-law scaling region for $\lambda=8$ is relatively unclear than the case of $\lambda=0$. It is possible that the finite-size effect on the correlation function becomes more significant for $\lambda=8$ and affects the power-law behavior. Indeed, we did see some hints of this behavior by comparing the results for $L=32$ and $40$ in Appendix F. There, we see that the power $p$ read from the correlation functions of different lattice sizes differs: the larger the lattice size $L$, we had the larger $p$. It is possible that larger $L$ gives a consistent estimate of $\eta$ with the other method. Furthermore, we observed that the correlation function is sensitive to the temperature, and thus it could be the source of ambiguity as well. Here we choose the peak temperature $\beta_{\mathrm{peak}}$ of the magnetic susceptibility, but the behavior of the correlation function changes even slightly by shifting the temperature. We examine the temperature dependence in Appendix G.
@@
All in all, we find reading critical exponents from the correlation function is less accurate than reading them from the finite size corrections of Binder parameters and magnetic susceptibility, both in the Heisenberg model and the local Heisenberg-dipolar model. This is in accord with what has been observed in the literature. We therefore report the values in Section V as our main results.
VII Summary and discussion
@@
Our obtained value of the critical exponent $\gamma$ turned out to be smaller than the estimation of the other method. There might exist a theoretical, as well as experimental, interpretation of this apparent smallness (if we assumed that it is indeed different from the dipolar fixed point value $\gamma_{\mathrm{dipolar}}$ [^6] The following interpretation was communicated to us by Slava Rychkov. We would like to thank him for his insight. . Recall that our simulation at finite $\lambda$ does not correspond to a true RG fixed point but rather to a crossover point, which has been pointed out by K. Ried et al. [35] (see also Ref. [36]). In such cases, it is argued that the effective susceptibility exponent $\gamma_{\mathrm{eff}}=\frac{d\log\chi}{d\log\beta}$ may show a significant dip away from the fixed point [5]. In our case, the deviation from the RG fixed point is mainly caused by finite $\lambda$, and our definition of $\gamma$ is different from $\gamma_{\mathrm{eff}}$, but it is plausible that the apparent smallness of our obtained $\gamma$ may be attributed to the similar effects. We also note that there is experimental evidence for such an effect in dipolar ferromagnets (see e.g. [37] and reference therein).
Therefore, one important future direction is to clarify the $\lambda$ dependence of critical exponents and estimate these values at the dipolar fixed point of Aharony and Fisher. An extrapolation focusing on $B=\langle(\nabla\cdot\vec{S})^{2}\rangle$ instead of $\lambda\rightarrow\infty$ as shown in Appendix  E might be a realistic approach in these future analyses.
Another important future problem is to consider the thermodynamic limit. As discussed, our results up to $L=40$ may still have a large finite-size effect. To do that, the improvement of simulation algorithms would be necessary. In this paper, we used the simplest Metropolis algorithm and encountered severe autocorrelation around the critical temperature, namely the critical slowing down, which prevented us from taking a larger lattice size than $L=40$, in particular with larger $\lambda$. For more precise determination of the critical exponents, the cluster algorithm [38, 39] can deal with this problem since it efficiently reduces the autocorrelation by a non-local update of the spin configuration. However, it is not straightforward to apply the algorithm to the local Heisenberg-dipolar model due to the next-nearest-neighbor interaction, the first term in the second line of Eq. (16). In this case, the generalized versions of the cluster algorithm will be helpful [40, 41]. Alternatively, it is also interesting to simulate the long-range dipolar interaction directly without using the local description. Indeed, the efficient algorithms for spin models with non-local interaction are proposed in Refs. [42, 43].
@@
Finally, we wish there would be a continuous effort to experimentally verify the dipolar effects in real ferromagnets  [^7] In the Chinese character, the root of magnetism is associated with a stone with benevolence. Why? In ancient Chinese stories, it is said that the way the ferromagnet attracts iron rocks makes it look as if an infant is naturally attracted to his or her mother. Probably the dipolar magnet will embody more benevolence because the magnetism from it is divergenceless, so it appears to be more gentle and embracing. So far in human history, the distinct features of dipolar magnetism have been observed in esoteric materials such as Europium compounds, but the universal nature of the benevolence should also be seen in more conventional materials such as iron or nickel in future experiments, just with more care..
@@ Appendix B Hubbard-Stratonovich transformation
Mapping of the parameters in the continuum field theory of Eq. (9) and a lattice model in Section III is non-trivial. Specifically, in Eq. (9), the term that imposes an important constraint in the local Heisenberg-dipolar model may acquire a multiplicative factor when mapped to the lattice model at finite $\lambda$. We derive the factor from the Hubbard-Stratonovich transformation within the mean field approximation.
@@
$1$ $=\quantity(\frac{\beta}{2\pi})^{|\Lambda|/2}\frac{1}{\sqrt{\det K}}$
$\quad\times\int\quantity(\prod_{l\in\Lambda}d\phi_{l})\exp\quantity(-\frac{\beta}{2}\sum_{n,m\in\Lambda}\vec{\phi}_{n}\cdot K^{-1}_{nm}\vec{\phi}_{m}),$ (46)
@@
$\quad\quad\left.\left.-\sum_{n\in\Lambda}2(\vec{\phi}_{n}+\vec{h}_{n})\cdot\vec{S}_{n}\right]\right\},$ (47)
@@
$Z$ $\sim\prod_{n=1}^{|\Lambda|}\int_{-\infty}^{\infty}d\phi_{n}\exp\left\{-\frac{\beta}{2}\sum_{n,m}\vec{\phi}_{n}\cdot(K^{-1})_{nm}\vec{\phi}_{m}\right$.
$\quad\quad\left.+\frac{\pi\beta^{2}\mu^{2}}{3}\sum_{n}(\vec{\phi}_{n}+\vec{h}_{n})^{2}\right\}.$ (48)
@@
$Z$ $\sim\int\mathcal{D}\vec{\phi}\exp\left(-\frac{\beta}{2}\int_{x,y}\vec{\phi}(x)\cdot K^{-1}(x-y)\vec{\phi}(y)\right$.
$\quad\quad\left.+\frac{\pi\beta^{2}\mu^{2}}{3}\int_{x}(\vec{\phi}(x)+\vec{h})^{2}\right),$ (49)
@@
[…]
@@
$\int d^{3}xd^{3}y\ \vec{\phi}(x)\cdot\tilde{K}(x-y)\vec{\phi}(y)$
$=\int d^{3}x\quantity(-J\theta\phi_{i}\phi_{i}+Ja^{2}(\partial_{i}\phi_{j})^{2}-2a^{2}\lambda(\partial_{i}\phi_{i})^{2}),$ (51)
@@
$\int d^{3}xd^{3}y\ \vec{\phi}(x)K^{-1}(x-y)\vec{\phi}(y)$
$\simeq\int d^{3}x\quantity(-\frac{1}{J\theta}\phi_{i}\phi_{i}-\frac{a^{2}}{J\theta^{2}}(\partial_{i}\phi_{j})^{2}+\frac{2a^{2}\lambda}{J^{2}\theta^{2}}(\partial_{i}\phi_{i})^{2}).$ (52)
@@
Consequently, we show that the normalization of $\lambda$ here is the one introduced in section II when we take J=1J=-1 in the lattice Hamiltonian (14) as a corresponding model.
@@ Appendix C Determination of the peak position of $\chi_{m}$
$\beta_{\mathrm{peak}}^{(L=32)}$ $0.689135(17)$ $0.353046(8)$ $0.320598(12)$
$\beta_{\mathrm{peak}}^{(L=40)}$ $0.690190(12)$ $0.352883(11)$ $0.320374(6)$
$\beta_{c}$ $0.693035(11)(^{+18}_{-20})$ - $0.319844(18)(^{+43}_{-41})$
@@ Appendix D Fitting results of the interpolating functions
Here, we summarize the best-fit values of the parameters $\{c_{i}(L)\}$ and $\{d_{i}(L)\}$ of the interpolating functions (30) and (36) which are obtained in Sections V.2 and V.5 in Table 5. The fitting errors of these values are large in some cases because the $\chi^{2}$-fit suffers from a relatively flat direction in the parameter spaces of $\{c_{i}(L)\}$ and $\{d_{i}(L)\}$. Even though the fitting parameters have large errors, the interpolated values have smaller errors, used to obtain the intersection and gradient, etc. Thus, it does not cause a problem in our analyses.
@@
Table 5: The best-fit values of the parameters $\{c_{i}(L)\}$ and $\{d_{i}(L)\}$ in the interpolating function (30) and (36) for $\lambda=0$ (top) and 88 (bottom). These values are normalized by $L^{2}$ or $L^{4}{10}^{3}$ to be almost the same order in these tables.
Appendix E Scaling Behavior in the $\lambda\to\infty$ limit
For the Heisenberg-dipolar model, the dipolar constraint is completely imposed in the $\lambda\to\infty$ limit. In this appendix, we test the validity of this extrapolation. As we measured the correlation function at $\lambda=0$, $4$, and $8$, it is feasible to perform the three-point extrapolation to $\lambda\to\infty$, assuming no phase transition at the finite $\lambda$. However, a naive extrapolation by a power-function of $1/\lambda$ cannot apply to $\lambda=0$. Therefore, we use $B=\sum_{n}B_{n}/V$ instead of $\lambda$ as a parameter of the dipolar constraint, where $B_{n}=\expectationvalue*{(\vec{\nabla}\cdot\vec{S}_{n})^{2}}$ is defined in Section IV.1. Since $B$ is finite at $\lambda=0$ and vanishes in the $\lambda\to\infty$ limit, the extrapolation to $B\to 0$ is well-defined.
In Figure 16, the critical exponent $\eta$ obtained from the correlation function is plotted against $B$ for the three cases of $\lambda$, where $\eta=p-1$ is taken from Table 3 for each element of the correlation function at $L=40$. The data points show almost linear behavior, implying $\eta$ and $B$ obey the same scaling for $\lambda$. Therefore, we fit the data points by a linear function, $\eta=\eta_{0}+aB$, and obtain $(\eta_{0},a)=(-0.49(4),0.37(5))$, (0.81(2),0.54(4))(-0.81(2),0.54(4)), and (0.84(2),0.62(4))(-0.84(2),0.62(4)) for the $xx$, yyyy, and $zz$ elements, respectively. Note that the extrapolated value $\eta_{0}$ is the result at the finite lattice size $L$ and is sensitive to $L$. Furthermore, as pointed out in Refs. [35, 36, 37], the $\lambda$ dependence of the exponents might be nonmonotonic. Increasing the number of data points for $\lambda$ would reveal the entire behavior of this model.
Figure 16: The critical exponent $\eta$ obtained from each diagonal element of the correlation function is plotted against $B=\sum_{n}B_{n}/V$ for the lattice size $L=40$. The symbols of circle-blue, square-green, and diamond-red correspond to the result of the $xx$, yyyy, and $zz$ elements, respectively. The horizontal error bar of $B$ is invisibly small. The fitting lines of $\eta=\eta_{0}+aB$ and the extrapolated values are depicted by the solid lines and plus symbols, respectively.
Appendix F Finite size effect on the correlation function
Here, we discuss the system-size dependence of the correlation function. In Figure 17, the correlation functions at the different system sizes $L=32$ and $40$ are compared, where the forward-backward average $C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$ is taken. The square-red data denote the result at the peak temperatures $\beta_{\mathrm{peak}}$ of the magnetic susceptibility on $L=40$ lattice, previously shown in Section VI, The circle-blue data are on $L=32$ lattice, where we tune the value of $\beta$ as $\beta_{\mathrm{peak}}=0.689135$ and $0.320598$ for $\lambda=0$ and $8$, respectively. We also depict the fitting lines of $c/|\vec{x}|^{p}$ in Figure 17, where the best-fit values are shown in Table 6 and 3 for $L=32$ and $40$, respectively.
Figure 17: The forward-backward averages $C_{ij}(\vec{x})+C_{ij}(-\vec{x})]/2$ of the diagonal elements of the correlation function are plotted against $|\vec{x}|$ in log-log scale for $\lambda=0$ (top) and 88 (bottom). The left, center, and right columns show the $xx$, yyyy, and $zz$ elements, where the circle-blue and diamond-red symbols correspond to $L=32$ and $40$, respectively. The solid lines depict the fitting results by $c/|\vec{x}|^{p}$.